Evaluvate (-27 1/9) - (-54 1/3)
Answer by edjones(8007)

(-27 1/9) - (-54 1/3)
=(-27 1/9) + (54 1/3)
=-244/9 + 163/3
=-244/9 + 489/9
=245/9
.
Ed